OpenStudy (anonymous):

Solve the system 3x – 2y = 7 and 9x – 6y = 3 What is the solution to the system?

OpenStudy (jamesj):

If you divide the second equation throughout by 3 we get 3x - 2y = 1 Now, is that consistent with the first equation, 3x - 2y = 7 ?

OpenStudy (anonymous):

Umm yes?

OpenStudy (jamesj):

No. Because the second equation implies 3x - 2y = 1 But the first equation is 3x - 2y = 7 If both equations where true, then that would imply that 7 = 1 which cannot be true!! We say that this system of equations is inconsistent, which means it leads to nonsense. Hence there are no solutions to both equations.

OpenStudy (jamesj):

Make sense?

OpenStudy (anonymous):

YES!! Thank you for explaining

OpenStudy (jamesj):

Great